I would like someone to show me how to solve this equation,(3/x+3)-(1/x-2)=(5/2x+6)
Thank you

Answer by lwsshak3(11628)

solve this equation:
(3/x+3)-(1/x-2)=(5/2x+6)
3/(x+3)-1/(x-2)=5/2(x+3)
LCD:2(x-2)(x+3)
3*2(x-2)-2(x+3)=5(x-2)
6x-12-2x-6=5x-10
x=-8